Sunland Company provides for bad debt expense at the rate of 2% of accounts receivable. The following data are available for 2018:

Allowance for doubtful accounts, 1/1/18 (Cr.) $ 13100
Accounts written off as uncollectible during 2018 9600
Ending accounts receivable 1198000


The Allowance for Doubtful Accounts balance at December 31, 2018, should be

2- Crane Company purchased equipment and these costs were incurred:

Cash price $65400
Sales taxes 4100
Insurance during transit 660
Installation and testing 830
Total costs

$70990


Crane will record the acquisition cost of the equipment as
Answer #1
1
Allowance for Doubtful Accounts balance at December 31, 2018=1198000*2%= $23960
2
Cost of the equipment =65400+4100+660+830=$70990